# Signed Networks: theory, methods, and applications

### v1

@@ 4.3 Sign prediction
Link prediction [kumarLinkPredictionTechniques2020] refers to the task of inferring missing edges in a network from its observed structure. In unsigned networks, this problem has received extensive attention due to its wide range of applications, like forecasting future interactions between agents and reconstructing incomplete datasets [cimini2021reconstructing]. When extending link prediction to signed networks, the problem splits into two related but distinct tasks: determining whether an edge exists between two nodes, and, assuming it exists, predicting its sign. In this book, we focus exclusively on the latter, as link prediction is often not signed-network-specific. Existing approaches to sign prediction can be grouped into four main categories. The first relies on structural balance, using measures of imbalance to infer the most aligned sign assignment. The second treats sign prediction as a classification problem, where edges are represented through structural features and their signs are predicted with statistical learning models. A third family uses matrix factorisation or completion, exploiting low-rank structure in the adjacency matrix to recover missing signs. Finally, deep learning methods learn edge representations directly from the graph, often through neural architectures that capture non-linear patterns. In the following, we describe a few methods from each category, and a comprehensive list can be found in Table LABEL:tab:sign-prediction.
Methods based on balance theory attribute a sign to an edge $(i,j)$ according to its consistency with the structural balance in the rest of the network [guhaPropagationTrustDistrust2004]. In this framework, the product of edge signs along a path encodes indirect influence: an even number of negative edges favours a positive relation, while an odd number favours a negative one. To aggregate these influences, one considers all $i\to j$ walks, typically under-weighting longer walks. A common formalisation uses powers of the signed adjacency matrix with a penalisation factor $c_{k}$, akin to the walk-based balance indices discussed in Section 3.5:
@@
Here, $\Gamma_{ij}$ measures the (weighted) difference between the number of positive and negative walks between $i$ and $j$; therefore, $\operatorname{sgn}(\Gamma_{ij})$ can be used as the edge sign prediction, with $|\Gamma_{ij}|$ providing a confidence score. Several penalisation schemes have been explored, such as exponential decay $c_{k}=\beta^{k}$ [guhaPropagationTrustDistrust2004, chiangExploitingLongerCycles2011, singhMeasuringBalanceSigned2017, kirkleyBalanceSignedNetworks2019] and factorial decay $c_{k}=\tfrac{1}{k!}$ [kunegisSlashdotZooMining2009, diazdiazGremban]. Beyond adjacency-based formulations, alternatives like the effective resistance and diffusion kernel have also been proposed [kunegisSlashdotZooMining2009, kunegisSpectralAnalysisSigned2010]. Another line of work treats positive and negative edges separately; for example, Guha et al. [guhaPropagationTrustDistrust2004] reported that the best results arise when the negative influence propagates at most once, leading to expressions such as $\Gamma=(A^{+})^{k}A^{-}$. Despite these variations, the underlying logic is always that the local patterns of balanced and unbalanced relations act as “votes” on the most consistent sign between two nodes.
@@
The second category of sign prediction methods frames it as a classification problem, where each edge is described by a vector of structural features that are then passed to a supervised learning algorithm. The pioneering work of Leskovec et al. [leskovecPredictingPositiveNegative2010] introduced this methodology using logistic regression. They proposed two main families of edge features. The first are degree-based features, which include the positive and negative in- and out-degrees of each endpoint, together with the edge embeddedness (defined as the number of common neighbours shared by the two endpoints). The second are triad-based features, obtained by counting how many triads of each type include the edge under consideration. Logistic regression then uses these features to predict the edge sign. This approach yields interpretable models, since the learned features’ weights can be directly related to balance and status theories. Its predictive power, however, depends strongly on embeddedness: triad features become informative only when the endpoints share many neighbours, so performance deteriorates for edges with low embeddedness. Moreover, because it relies exclusively on degree- and triad-based features, the method cannot capture balance patterns that emerge through longer cycles.
@@
To address these limitations, Chiang et al. [chiangExploitingLongerCycles2011] extended the feature construction so that the representation of an edge $(i,j)$ includes counts of walks forming cycles of length up to $k$. For example, to capture the influence of 4-cycles, they considered terms such as $(A^{+})^{4}_{ij}$, $(A^{+})^{3}_{ij}(A^{-})_{ij}$, $(A^{+})^{2}_{ij}(A^{-})^{2}_{ij}$, $(A^{+})_{ij}(A^{-})^{3}_{ij}$, and $(A^{-})^{4}_{ij}$. Incorporating these higher-order cycle features into logistic regression significantly improved predictive performance, particularly for edges with low embeddedness where triadic features provide no information. Building on the idea of going beyond triads, Papaoikonomou et al. [papaoikonomouPredictingEdgeSigns2014] proposed to represent edges through frequent subgraph patterns (i.e., motifs) in their local ego-graphs, including them as a feature in the support vector machine model used for classification. Their results showed that these motifs carry strong discriminative power. More recently, Derr et al. [derrBalanceSignedBipartite2019] addressed the case of bipartite signed networks, which by definition cannot have triangles or any loop of odd length. They replaced triad-based features with bipartite motifs such as “caterpillars” and “butterflies”  [^18] Signed butterflies are cycles of length 4, and caterpillars are paths of length 3 that can become a butterfly by adding one edge., and used them in a classifier to predict edge signs.
@@
The third category for sign prediction in signed networks treats the adjacency matrix $A$ as a partially observed matrix and approximate it by a low-rank factorization $A\approx UV^{\top}$, where $U\in\mathbb{R}^{n\times d}$ and $V\in\mathbb{R}^{n\times d}$ ($d\ll n$) [forsati2015pushtrust, tang2016recommendations, derrBalanceSignedBipartite2019]. The factorization is attained by minimizing a loss $\ell$ over the observed entries, for example [forsati2015pushtrust]
@@
Where $\Omega$ is the set of known signs and $\ell$ encourages $u_{i}^{\top}v_{j}$ to be positive for friendly edges and negative for antagonistic ones. $\|U\|_{F}^{2}+\|V\|_{F}^{2}$ is a regularisation term that prevents overfitting by penalising excessively large embeddings, and whose importance is determined by the parameter $\lambda$. After optimisation, the sign of an unobserved edge is predicted by $\hat{A}_{ij}=\operatorname{sign}(u_{i}^{\top}v_{j}),$ where $u_{i}$ (the $i$-th row of $U$) and $v_{j}$ (the $j$-th row of $V$) represent the latent vectors of nodes $i$ and $j$. Alternatively, instead of assigning a fixed $\pm 1$ value to each interaction, probabilistic variants weight edges through a logistic or hinge-based function. A related method, introduced by Hsieh et al. [hsiehLowRankModeling2012], builds on the observation that the adjacency matrix of weakly balanced complete networks has rank equal to the number of factions. This motivates formulating sign prediction as a matrix completion problem: given a partially observed $A$, the task is to recover the missing entries under a low-rank constraint. In practice, because the rank function is non-convex, it is typically replaced by the convex trace norm. The resulting matrix is used to predict the missing signs. Similarly, Ye et al. [ye_predicting_2013] introduced a matrix-based approach with transfer learning. In their framework, two networks (the source and the target) are embedded into a shared latent space. Structural patterns are first learned from the well-labelled source and then used to guide sign prediction in a sparsely labelled target.
Finally, the last category of techniques for sign prediction is based on deep learning approaches. Ruiz-Garcia et al [ruiz-garciaTriadicInfluenceProxy2023] used neural networks to predict the sign of social ties among teenagers from a measure of imbalance, the triadic influence. Interestingly, adding node-level features such as gender or prosociality did not improve prediction, meaning that the structure of connections between individuals provides more information about their specific characteristics. Alternatively, GNN-based models have been adapted to message passing to handle both positive and negative edges. For instance, SELO encodes subgraphs around each link to learn edge embeddings [fangSignedSubgraphEncoding2023], SIHG places embeddings in hyperbolic space with attention mechanisms that separate friends and foes [luoInterpretableSignedLink2021], and Chen et al. introduce graph augmentation strategies to filter unreliable negative edges and improve robustness [chenBalancingAugmentationEdgeUtility2023].
@@ 4.4 Community detection
Community detection aims to detect mesoscopic modular structures in large, complex systems; this is, to find groups of nodes in a network that share common characteristics [fortunatoCommunityDetectionGraphs2010]. This task has attracted considerable attention in the scientific literature, mainly for two reasons: first, because communities reduce the complexity of a network and let us focus on interpretable modules that reflect the underlying organisation, and second, because of the algorithmic challenges of designing a robust and efficient community detection method. In social networks, these modules might correspond to interest groups, political factions, or geographic clusters. In scientific collaboration networks, they are typically identified with scientific fields and research topics [girvanCommunityStructureSocial2002]. In brain networks, they often align with functionally specialised regions involved in specific cognitive tasks [spornsModularBrainNetworks2016]. In ecological and metabolic networks, communities can reveal trophic levels or biochemical pathways [gauzensTrophicGroupsModules2015].
@@
Most methods for community detection in unsigned networks build on the idea that nodes within the same group tend to be more densely connected than nodes in different groups [fortunatoCommunityDetectionGraphs2010]. For instance, when considering face-to-face contacts, a person is more likely to interact with coworkers than with people working in another location. However, this structural view alone may not always capture meaningful community structure, especially when additional information, such as node or edge attributes, is available. These attributes can uncover groupings that topology alone might miss.
@@
This limitation becomes evident in signed networks, where the information carried in the edge sign is relevant to detecting and interpreting hostile communities. This leads to a new conceptualisation of a signed network community as a group of nodes with internal agreement and external disagreement [^19] To distinguish between these two conceptualisations of community, the term faction is often used to refer to communities in signed networks.. In this perspective, a good partition minimises the number of negative edges within communities and the number of positive edges between them [^20] Note that this objective is equivalent to minimising the frustration of a given partition, a topic discussed in Section 3.5. This definition of community is closely connected to structural balance theory: a network with a clear signed community structure typically exhibits a high level of structural balance. In the limit of perfectly balanced networks, the community structure becomes trivial: the network can be partitioned into mutually antagonistic balanced factions, characterised by only positive links within each faction and only negative links between them. Figure 10 shows the difference between community detection in unsigned networks, where the goal is typically to maximise within-group edge density, and signed networks, where the goal is typically frustration minimisation. It is important to recognise that community detection always relies on assumptions about the type of structure we expect, and different assumptions can reveal different communities with different interpretations [peelGroundTruth].
@@
Finding communities by minimising frustration is typically an NP-hard problem, which makes exact solutions computationally infeasible for large networks. As a result, most approaches rely on approximations. Several algorithms originally developed for unsigned networks have been generalised to signed ones, such as Infomap [esmailianCommunityDetectionSigned2015], walk-based strategies [suAlgorithmBasedPositive2017], and matrix completion methods [hsiehLowRankModeling2012]. Broadly, most methods proposed in the literature can be grouped into three main families: heuristic algorithms that optimise a quality function, spectral methods that exploit the spectrum of the Laplacian or adjacency matrix, and approaches based on statistical inference. Tables A1, A2, A3 and A4 show a comprehensive view of the methods analysed in this book. In the remainder of this section, we outline the most popular strategies and representative methods for community detection in signed networks. For comprehensive comparisons and further methodological details, we refer the reader to [tangSurveySignedNetwork2016, chiangPredictionClusteringSigned2013a, tomassoAdvancesScalingCommunity2022, traagPartitioningSignedNetworks2018, schaubStructuredNetworksCoarsegrained2019, candellone2025community].
@@ 4.4.1 Quality function optimization
We first explore the most important quality functions proposed to detect communities in signed networks. The first one was introduced by Doreian and Mrvar [doreianPartitioningApproachStructural1996]. Their formulation combines two criteria: the number of positive links between different communities $m^{+}(C_{a},C_{b})$, i.e., the out-group agreement, and the number of negative links within each community $m^{-}(C_{a})$, i.e., the in-group disagreement. The objective function to minimise is defined as a weighted average of these two quantities:
@@
where $0\leq\alpha\leq 1$ controls the relative importance of the two terms. Note that, for $\alpha=0.5$, this quality function reduces to the frustration index of a signed network [hararyMeasurementStructuralBalance1959]. An illustration of the algorithmic procedure to find the partition of the network via frustration minimisation is depicted in Figure 11. The authors later extended this approach to bipartite (two-mode) networks, adapting the quality function to account for the structural properties specific to such systems [mrvarPartitioningSignedTwoMode2009], and also to detect structures consistent with relaxed structural balance [doreianPartitioningSignedSocial2009] (see Section 3.4).
A conceptually different approach to quality-function optimisation was later introduced through extensions of modularity for signed networks [newmanFindingEvaluatingCommunity2004]. The modularity function compares the observed edge density within communities to that expected under a null model in which connections are randomised but node degrees are preserved. Formally, given a partition $C$ of the node set, the modularity is defined as
@@
To extend this idea to signed networks, Gómez, Jensen, and Arenas [gomezAnalysisCommunityStructure2009] proposed treating positive and negative links separately, each with its own null model. The signed modularity is defined as
@@
A more general framework was proposed by Traag and Bruggeman [traagCommunityDetectionNetworks2009], rooted in statistical mechanics. Extending the work of Reichardt and Bornhold for unsigned networks [reichardtStatisticalMechanicsCommunity2006], they interpret the quality function as the Hamiltonian of a Potts model that separates the contribution of positive and negative edges in the same spirit as Gómez et al. [gomezAnalysisCommunityStructure2009].
@@
where $A_{ij}$ is the signed adjacency matrix, $p^{+}_{ij}$ and $p^{-}_{ij}$ are the expected positive and negative weights between nodes $i$ and $j$ under the null model, and $\gamma^{+},\gamma^{-}$ are resolution parameters controlling the relative contribution of each term. Typically, the chosen null model is the configuration model, so that $p_{ij}^{\pm}=\frac{k_{i}^{\pm}k_{j}^{\pm}}{2m^{\pm}}$. Under this choice, and setting $\gamma^{+}=\gamma^{-}=1$, the Hamiltonian reduces to the modularity function from [gomezAnalysisCommunityStructure2009]. Alternatively, choosing $\gamma^{+}=\gamma^{-}$ and $\alpha=\frac{1}{2}$, the Hamiltonian becomes the frustration index (Eq. 5). This shows that the quality function in Eq. 6 unifies and generalises both frustration-based and modularity-based formulations into a single quality function. Moreover, the interpretation of the quality function as a Hamiltonian makes it ideally suited for use in conjunction with simulated annealing minimisation algorithms. Finally, the parameters $\gamma^{\pm}$ can be interpreted as resolution parameters that circumvent the infamous resolution limit of the modularity function [fortunatoResolutionLimitCommunity2007]. The main disadvantage of the method is the necessity to tune those two additional parameters. Some subsequent works have proposed ways to choose these parameters in a principled way (see [esmailianCommunityDetectionSigned2015]).
@@
It is important to note that quality function-based algorithms, especially those based on modularity, inherit the same limitations observed in the unsigned case [peixotoDescriptiveVsInferential2023], such as the resolution limit and the tendency to overfit the data [guimeraModularityFluctuationsRandom2004]. Despite these drawbacks, modularity remains a popular quality function for signed graphs (see e.g. [anchuriCommunitiesBalanceSigned2012] and [amelioCommunityMiningSigned2013]).
@@
Once a quality function is fixed, optimisation algorithms are used to find the best partition. Since optimising these functions is NP-hard [brandesModularityClustering2008], heuristic methods are typically required. Two of the most popular ones are the Louvain algorithm (and its extensions such as Leiden [traag2019louvain]) and simulated annealing. The Louvain algorithm  [blondelFastUnfoldingCommunities2008] is a greedy, hierarchical method that iteratively optimises modularity by locally reassigning nodes to communities, aggregating communities into super-nodes, and then repeating the process. In Louvain, nodes cannot move into communities of non-neighbouring nodes, a restriction that in signed networks often leads to fragmented partitions, particularly when negative edges are abundant. To solve this issue, Pougué-Biyong et al. have recently proposed SignedLouvain [pougue-biyongSignedLouvainLouvainSigned2024], a modification of the Louvain algorithm that allows nodes to explore communities two steps away via negative links. This targeted relaxation lets the algorithm recover balanced mesoscopic structures while remaining fast and consistent.
@@
Simulated annealing, on the other hand, offers a more flexible approach that is particularly effective when optimising global energy-like functions such as frustration [kirkpatrickOptimizationSimulatedAnnealing1983]. It allows for configurations that increase the energy of the systems during early phases, to escape poor local optima that hinder greedy algorithms. This makes it suitable for frustration-based formulations, which often exhibit a rugged energy landscape. Nonetheless, its computational cost is higher, and it lacks the fast multi-scale aggregation mechanism of Louvain. In practice, Louvain-type methods work when optimising modularity-like scores, while simulated annealing provides a robust, albeit slower, alternative for global criteria or when precise partitions are needed in small to medium-sized graphs. Other approaches to minimise quality functions include the Parallel Iterative Local Search algorithm [mendoncaRelevanceNegativeLinks2015], and integer programming approaches [arefModelingComputationalStudy2020, arefDetectingCoalitionsOptimally2020, arefBalanceFrustrationSigned2019]

### v2

@@ 4.3 Sign prediction
Link prediction [221] refers to the task of inferring missing edges in a network from its observed structure. In unsigned networks, this problem has received extensive attention due to its wide range of applications, like forecasting future interactions between agents and reconstructing incomplete datasets [73]. When extending link prediction to signed networks, the problem splits into two related but distinct tasks: determining whether an edge exists between two nodes, and, assuming it exists, predicting its sign. In this book, we focus exclusively on the latter, as link prediction is often not signed-network-specific. Existing approaches to sign prediction can be grouped into four main categories. The first relies on structural balance, using measures of imbalance to infer the most aligned sign assignment. The second treats sign prediction as a classification problem, where edges are represented through structural features and their signs are predicted with statistical learning models. A third family uses matrix factorisation or completion, exploiting low-rank structure in the adjacency matrix to recover missing signs. Finally, deep learning methods learn edge representations directly from the graph, often through neural architectures that capture non-linear patterns. In the following, we describe a few methods from each category, and a comprehensive list can be found in Table LABEL:tab:sign-prediction.
Methods based on balance theory attribute a sign to an edge $(i,j)$ according to its consistency with the structural balance in the rest of the network [157]. In this framework, the product of edge signs along a path encodes indirect influence: an even number of negative edges favours a positive relation, while an odd number favours a negative one. To aggregate these influences, one considers all $i\to j$ walks, typically under-weighting longer walks. A common formalisation uses powers of the signed adjacency matrix with a penalisation factor $c_{k}$, akin to the walk-based balance indices discussed in Section 3.5:
@@
Here, $\Gamma_{ij}$ measures the (weighted) difference between the number of positive and negative walks between $i$ and $j$; therefore, $\operatorname{sgn}(\Gamma_{ij})$ can be used as the edge sign prediction, with $|\Gamma_{ij}|$ providing a confidence score. Several penalisation schemes have been explored, such as exponential decay $c_{k}=\beta^{k}$ [157, 68, 356, 212] and factorial decay $c_{k}=\tfrac{1}{k!}$ [223, 92]. Beyond adjacency-based formulations, alternatives like the effective resistance and diffusion kernel have also been proposed [223, 224]. Another line of work treats positive and negative edges separately; for example, Guha et al. [157] reported that the best results arise when the negative influence propagates at most once, leading to expressions such as $\Gamma=(A^{+})^{k}A^{-}$. Despite these variations, the underlying logic is always that the local patterns of balanced and unbalanced relations act as “votes” on the most consistent sign between two nodes.
@@
The second category of sign prediction methods frames it as a classification problem, where each edge is described by a vector of structural features that are then passed to a supervised learning algorithm. The pioneering work of Leskovec et al. [231] introduced this methodology using logistic regression. They proposed two main families of edge features. The first are degree-based features, which include the positive and negative in- and out-degrees of each endpoint, together with the edge embeddedness (defined as the number of common neighbours shared by the two endpoints). The second are triad-based features, obtained by counting how many triads of each type include the edge under consideration. Logistic regression then uses these features to predict the edge sign. This approach yields interpretable models, since the learned features’ weights can be directly related to balance and status theories. Its predictive power, however, depends strongly on embeddedness: triad features become informative only when the endpoints share many neighbours, so performance deteriorates for edges with low embeddedness. Moreover, because it relies exclusively on degree- and triad-based features, the method cannot capture balance patterns that emerge through longer cycles.
@@
To address these limitations, Chiang et al. [68] extended the feature construction so that the representation of an edge $(i,j)$ includes counts of walks forming cycles of length up to $k$. For example, to capture the influence of 4-cycles, they considered terms such as $(A^{+})^{4}_{ij}$, $(A^{+})^{3}_{ij}(A^{-})_{ij}$, $(A^{+})^{2}_{ij}(A^{-})^{2}_{ij}$, $(A^{+})_{ij}(A^{-})^{3}_{ij}$, and $(A^{-})^{4}_{ij}$. Incorporating these higher-order cycle features into logistic regression significantly improved predictive performance, particularly for edges with low embeddedness where triadic features provide no information. Building on the idea of going beyond triads, Papaoikonomou et al. [303] proposed to represent edges through frequent subgraph patterns (i.e., motifs) in their local ego-graphs, including them as a feature in the support vector machine model used for classification. Their results showed that these motifs carry strong discriminative power. More recently, Derr et al. [85] addressed the case of bipartite signed networks, which by definition cannot have triangles or any loop of odd length. They replaced triad-based features with bipartite motifs such as “caterpillars” and “butterflies”  [^18] Signed butterflies are cycles of length 4, and caterpillars are paths of length 3 that can become a butterfly by adding one edge., and used them in a classifier to predict edge signs.
@@
The third category for sign prediction in signed networks treats the adjacency matrix $A$ as a partially observed matrix and approximate it by a low-rank factorization $A\approx UV^{\top}$, where $U\in\mathbb{R}^{n\times d}$ and $V\in\mathbb{R}^{n\times d}$ ($d\ll n$) [123, 378, 85]. The factorization is attained by minimizing a loss $\ell$ over the observed entries, for example [123]
@@
Where $\Omega$ is the set of known signs and $\ell$ encourages $u_{i}^{\top}v_{j}$ to be positive for friendly edges and negative for antagonistic ones. $\|U\|_{F}^{2}+\|V\|_{F}^{2}$ is a regularisation term that prevents overfitting by penalising excessively large embeddings, and whose importance is determined by the parameter $\lambda$. After optimisation, the sign of an unobserved edge is predicted by $\hat{A}_{ij}=\operatorname{sign}(u_{i}^{\top}v_{j}),$ where $u_{i}$ (the $i$-th row of $U$) and $v_{j}$ (the $j$-th row of $V$) represent the latent vectors of nodes $i$ and $j$. Alternatively, instead of assigning a fixed $\pm 1$ value to each interaction, probabilistic variants weight edges through a logistic or hinge-based function. A related method, introduced by Hsieh et al. [188], builds on the observation that the adjacency matrix of weakly balanced complete networks has rank equal to the number of factions. This motivates formulating sign prediction as a matrix completion problem: given a partially observed $A$, the task is to recover the missing entries under a low-rank constraint. In practice, because the rank function is non-convex, it is typically replaced by the convex trace norm. The resulting matrix is used to predict the missing signs. Similarly, Ye et al. [408] introduced a matrix-based approach with transfer learning. In their framework, two networks (the source and the target) are embedded into a shared latent space. Structural patterns are first learned from the well-labelled source and then used to guide sign prediction in a sparsely labelled target.
Finally, the last category of techniques for sign prediction is based on deep learning approaches. Ruiz-Garcia et al [332] used neural networks to predict the sign of social ties among teenagers from a measure of imbalance, the triadic influence. Interestingly, adding node-level features such as gender or prosociality did not improve prediction, meaning that the structure of connections between individuals provides more information about their specific characteristics. Alternatively, GNN-based models have been adapted to message passing to handle both positive and negative edges. For instance, SELO encodes subgraphs around each link to learn edge embeddings [117], SIHG places embeddings in hyperbolic space with attention mechanisms that separate friends and foes [251], and Chen et al. introduce graph augmentation strategies to filter unreliable negative edges and improve robustness [65].
@@ 4.4 Community detection
Community detection aims to detect mesoscopic modular structures in large, complex systems; this is, to find groups of nodes in a network that share common characteristics [125]. This task has attracted considerable attention in the scientific literature, mainly for two reasons: first, because communities reduce the complexity of a network and let us focus on interpretable modules that reflect the underlying organisation, and second, because of the algorithmic challenges of designing a robust and efficient community detection method. In social networks, these modules might correspond to interest groups, political factions, or geographic clusters. In scientific collaboration networks, they are typically identified with scientific fields and research topics [146]. In brain networks, they often align with functionally specialised regions involved in specific cognitive tasks [362]. In ecological and metabolic networks, communities can reveal trophic levels or biochemical pathways [143].
@@
Most methods for community detection in unsigned networks build on the idea that nodes within the same group tend to be more densely connected than nodes in different groups [125]. For instance, when considering face-to-face contacts, a person is more likely to interact with coworkers than with people working in another location. However, this structural view alone may not always capture meaningful community structure, especially when additional information, such as node or edge attributes, is available. These attributes can uncover groupings that topology alone might miss.
@@
This limitation becomes evident in signed networks, where the information carried in the edge sign is relevant to detecting and interpreting hostile communities. This leads to a new conceptualisation of a signed network community as a group of nodes with internal agreement and external disagreement [^19] To distinguish between these two conceptualisations of community, the term faction is often used to refer to communities in signed networks.. In this perspective, a good partition minimises the number of negative edges within communities and the number of positive edges between them [^20] Note that this objective is equivalent to minimising the frustration of a given partition, a topic discussed in Section 3.5. This definition of community is closely connected to structural balance theory: a network with a clear signed community structure typically exhibits a high level of structural balance. In the limit of perfectly balanced networks, the community structure becomes trivial: the network can be partitioned into mutually antagonistic balanced factions, characterised by only positive links within each faction and only negative links between them. Figure 10 shows the difference between community detection in unsigned networks, where the goal is typically to maximise within-group edge density, and signed networks, where the goal is typically frustration minimisation. It is important to recognise that community detection always relies on assumptions about the type of structure we expect, and different assumptions can reveal different communities with different interpretations [306].
@@
Finding communities by minimising frustration is typically an NP-hard problem, which makes exact solutions computationally infeasible for large networks. As a result, most approaches rely on approximations. Several algorithms originally developed for unsigned networks have been generalised to signed ones, such as Infomap [105], walk-based strategies [368], and matrix completion methods [188]. Broadly, most methods proposed in the literature can be grouped into three main families: heuristic algorithms that optimise a quality function, spectral methods that exploit the spectrum of the Laplacian or adjacency matrix, and approaches based on statistical inference. Tables A1, A2, A3 and A4 show a comprehensive view of the methods analysed in this book. In the remainder of this section, we outline the most popular strategies and representative methods for community detection in signed networks. For comprehensive comparisons and further methodological details, we refer the reader to [379, 67, 384, 390, 344, 54].
@@ 4.4.1 Quality function optimization
We first explore the most important quality functions proposed to detect communities in signed networks. The first one was introduced by Doreian and Mrvar [96]. Their formulation combines two criteria: the number of positive links between different communities $m^{+}(C_{a},C_{b})$, i.e., the out-group agreement, and the number of negative links within each community $m^{-}(C_{a})$, i.e., the in-group disagreement. The objective function to minimise is defined as a weighted average of these two quantities:
@@
where $0\leq\alpha\leq 1$ controls the relative importance of the two terms. Note that, for $\alpha=0.5$, this quality function reduces to the frustration index of a signed network [171]. An illustration of the algorithmic procedure to find the partition of the network via frustration minimisation is depicted in Figure 11. The authors later extended this approach to bipartite (two-mode) networks, adapting the quality function to account for the structural properties specific to such systems [280], and also to detect structures consistent with relaxed structural balance [97] (see Section 3.4).
A conceptually different approach to quality-function optimisation was later introduced through extensions of modularity for signed networks [293]. The modularity function compares the observed edge density within communities to that expected under a null model in which connections are randomised but node degrees are preserved. Formally, given a partition $C$ of the node set, the modularity is defined as
@@
To extend this idea to signed networks, Gómez, Jensen, and Arenas [147] proposed treating positive and negative links separately, each with its own null model. The signed modularity is defined as
@@
A more general framework was proposed by Traag and Bruggeman [386], rooted in statistical mechanics. Extending the work of Reichardt and Bornhold for unsigned networks [325], they interpret the quality function as the Hamiltonian of a Potts model that separates the contribution of positive and negative edges in the same spirit as Gómez et al. [147].
@@
where $A_{ij}$ is the signed adjacency matrix, $p^{+}_{ij}$ and $p^{-}_{ij}$ are the expected positive and negative weights between nodes $i$ and $j$ under the null model, and $\gamma^{+},\gamma^{-}$ are resolution parameters controlling the relative contribution of each term. Typically, the chosen null model is the configuration model, so that $p_{ij}^{\pm}=\frac{k_{i}^{\pm}k_{j}^{\pm}}{2m^{\pm}}$. Under this choice, and setting $\gamma^{+}=\gamma^{-}=1$, the Hamiltonian reduces to the modularity function from [147]. Alternatively, choosing $\gamma^{+}=\gamma^{-}$ and $\alpha=\frac{1}{2}$, the Hamiltonian becomes the frustration index (Eq. 5). This shows that the quality function in Eq. 6 unifies and generalises both frustration-based and modularity-based formulations into a single quality function. Moreover, the interpretation of the quality function as a Hamiltonian makes it ideally suited for use in conjunction with simulated annealing minimisation algorithms. Finally, the parameters $\gamma^{\pm}$ can be interpreted as resolution parameters that circumvent the infamous resolution limit of the modularity function [124]. The main disadvantage of the method is the necessity to tune those two additional parameters. Some subsequent works have proposed ways to choose these parameters in a principled way (see [105]).
@@
It is important to note that quality function-based algorithms, especially those based on modularity, inherit the same limitations observed in the unsigned case [310], such as the resolution limit and the tendency to overfit the data [158]. Despite these drawbacks, modularity remains a popular quality function for signed graphs (see e.g. [13] and [11]).
@@
Once a quality function is fixed, optimisation algorithms are used to find the best partition. Since optimising these functions is NP-hard [49], heuristic methods are typically required. Two of the most popular ones are the Louvain algorithm (and its extensions such as Leiden [387]) and simulated annealing. The Louvain algorithm  [40] is a greedy, hierarchical method that iteratively optimises modularity by locally reassigning nodes to communities, aggregating communities into super-nodes, and then repeating the process. In Louvain, nodes cannot move into communities of non-neighbouring nodes, a restriction that in signed networks often leads to fragmented partitions, particularly when negative edges are abundant. To solve this issue, Pougué-Biyong et al. have recently proposed SignedLouvain [315], a modification of the Louvain algorithm that allows nodes to explore communities two steps away via negative links. This targeted relaxation lets the algorithm recover balanced mesoscopic structures while remaining fast and consistent.
@@
Simulated annealing, on the other hand, offers a more flexible approach that is particularly effective when optimising global energy-like functions such as frustration [213]. It allows for configurations that increase the energy of the systems during early phases, to escape poor local optima that hinder greedy algorithms. This makes it suitable for frustration-based formulations, which often exhibit a rugged energy landscape. Nonetheless, its computational cost is higher, and it lacks the fast multi-scale aggregation mechanism of Louvain. In practice, Louvain-type methods work when optimising modularity-like scores, while simulated annealing provides a robust, albeit slower, alternative for global criteria or when precise partitions are needed in small to medium-sized graphs. Other approaches to minimise quality functions include the Parallel Iterative Local Search algorithm [267], and integer programming approaches [19, 20, 22]